# DyFFPAD: Dynamic Fusion of Convolutional and Handcrafted Features for Fingerprint Presentation Attack Detection

### v2

@@ 3 Proposed Work
In this paper, an ensemble of deep and handcrafted features is proposed for the detection of spoofs in intra-sensor same-material and cross-material protocols of FPAD. It consists of two modules that collaborate to detect the PAs. The first module is a DNN that is fed with the features extracted by Local Phase Quantization (LPQ) and existing handcrafted features. The second module is the DenseNet CNN classifier which has shown remarkable performance while being validated on various image databases including imagenet [6]. Both of the modules are fused dynamically which results in better learning of them as compared with their individual performances. The details of LPQ, handcrafted features, DenseNet, and DyFFPAD are given in the following subsections.
@@ 3.1 Local Phase Quantization (LPQ)
LPQ [11] works on the Fourier transform’s blur-insensitive property. This makes LPQ robust against the redundant information and blur present in the input sample. We have utilized it as a prominent feature due to its capability of exploiting the minute information that gets missing in the fabrication process of a spoof. The formulation of the LPQ descriptor is denoted as Eq. 1.
@@ 3.2 Handcrafted Features
The fabrication process impacts the quality of the spoofs fabricated with various materials. According to recent research [32] [31], handcrafted features are proven to be useful while detecting PAs. In this work, we have utilized some existing handcrafted features [32], [44] including ridge-valley clarity, gabor quality, number of abnormal ridges and valleys, ridge-valley smoothness, frequency domain analysis, and orientation certainty level to estimate the quality of the input fingerprint sample. The details of these features is provided in the following subsections.
@@ 3.2.1 Ridge and Valley Smoothness (RVS and VWS)
It denotes the smoothness of ridge width and valley width which is exhibited by the live sample but not by the spoof. This irregularity in the spoofs is caused because of the varying elasticity levels of spoof fabrication materials. These materials do not possess the elasticity same as human skin. In addition to this, the pressure applied on the fingertip at the time of sample collection is also a potential reason behind this. RWS and VWS are computed block-wise by cropping the rotated block with a vertical ridge-valley structure followed by the binarization. Thereafter pixels of the block are labeled as ridge or valley with the help of a linear regression algorithm. After this, ridge and valley width is computed for each horizontal line of the block. Averaging the standard deviation of ridge and valley width is done to obtain RWS and VWS feature.
@@ 3.2.2 Frequency Domain Analysis (FDA)
FDA of a local block is extracted by computing the ridge-valley structure’s 1D signature. The frequency of the sinusoidal ridge-valley structure is calculated using the discrete Fourier transform of this 1D signature. Spoof fingerprint samples exhibit a varying frequency of sinusoidal ridge-valley patterns while it is consistent in live fingerprints.
@@ 3.2.4 Gabor Quality (G)
A filter bank with several orientations is applied to each pixel to compute gabor quality at the local level. For one or a few filters with orientations comparable to the block orientation, the Gabor response for a fingerprint block with a normal ridge-valley pattern remains high, but it is low and steady for the same block with an improper ridge-valley structure. Finally, the Gabor quality (G) of a block is calculated by taking the standard deviation of the output of the gabor filter bank.
@@ 3.2.5 Ridge-Valley Clarity (RVC)
The ridge and valleys present in a block of live fingerprints are found to be separated consistently. This happens due to the uniform elasticity property of human finger skin.This separation, on the other hand, differs in spoof samples due to inappropriate widths of ridges and valleys induced by the different elasticity levels of spoofing materials employed in spoof fabrication. The RVC feature is computed by dividing the original sample into fixed-sized blocks and then computing the average ridge and valley widths in the block.
3.2.6 Number of abnormal ridges and valleys ($R_{ab}$ and $V_{ab}$)
The width of ridges and valleys is found to be in a range of 5 to 10 pixels in samples of human fingers captured from a sensing device with a resolution of 500 dpi. Since the human skin and spoofing materials have different levels of elasticity, the width of ridges and valleys differs in corresponding samples. A ridge or valley in a local block is regarded as abnormal if its width divergence in various rows of the block exceeds a particular threshold.
@@ 3.2.7 Feature vector from quality features
The final feature vector which is represented as Eq. (2) is generated using the mean ($\mu$) and standard deviation ($\sigma$) of the features discussed above.
@@ 3.3 DenseNet-121
DenseNet[15] consists of convolutional layers connected to their previous layers in a feed-forward manner. The majority of CNN designs suffer from the vanishing gradient problem, which occurs when the Rectified Linear Unit (ReLU) is used as an activation function. In deeper CNNs, the gradient of the loss function approaches zero due to the use of ReLU activation functions. This phenomenon further impacts the learning process which results in overfitting of the model. As fingerprint images have limited colour and texture information, typical CNNs encounter this issue when they are utilized in this domain. This problem of vanishing gradient is solved by DenseNet which has skip connections among the layers inside convolutional blocks as shown in Fig. 2.
@@
DenseNet-121 comprises four dense blocks containing six, twelve, twenty-four, and sixteen convolutional layers, respectively. The internal architecture of DenseNet is illustrated in Figure 2. Each dense block performs three operations: batch normalization, ReLU activation and convolution, and aggregates feature maps from all its preceding layers. The output of a dense block is formulated as shown in Eq. (3).
@@
where $A_{0},A_{1},A_{n-1}$ represents the concatenation of all feature maps from the previous layers and $F_{n}$ denotes a function that performs batch normalization followed by a convolution operation. In addition, after each dense block, a transition block is utilized to execute a convolution operation with a kernel size of $1\times 1$. Following this convolution procedure, a pooling operator minimizes the size of feature maps after each dense block. Figure 3 depicts the internal architecture of the dense layer and transition layer.
@@ 3.4 Deep Neural Network (DNN)
DNN is a type of artificial neural network that has several fully connected layers between the input and output layers. The working of DNNs is inspired by the human brain as its neurons resemble similar characteristics as neurons of the brain. We have used DNN to test the individual performance of the handcrafted features and LPQ image descriptor and to design the proposed architecture.
@@ 3.5 Proposed DyFFPAD Model
The architecture of the proposed DyFFPAD is depicted in Fig. 4. It is a dynamic ensemble of handcrafted and convolutional features. The DNN works on the feature vectors generated from handcrafted features and LPQ image descriptors while the convolution base of DenseNet extracts convolutional features from the input samples. The Features extracted from both modules are combined to form a feature vector which is further fed as an input to a DNN that consists of fully connected layers with 48, 16, and 1 neuron respectively. In the forward pass of the training process, a confidence score is produced by the model. The loss of the output score with respect to the expected output is calculated which is back-propagated to the entire model including CNN as well as DNNs for the learning of their parameters. The proposed model works in an end-to-end manner for the forward and backward pass of the training process and reduces the complexities of the ensemble of handcrafted features and convolutional features for the accomplishment of a classification task.
@@ 3.5.1 Pre-processing of input samples
The fingerprint samples captured with the sensing devices have white space around the fingertip impression. This white space is meaningless and is required to be removed for the extraction of the desired features from fingerprint samples. The input fingerprint undergoes the pre-processing operation for the extraction of the region of interest from the input fingerprint sample.
@@ 3.5.3 Training of DyFFPAD
The proposed DyFFPAD model is trained from scratch on benchmark LivDet databases. To expedite the training process, we initialized the parameters of DenseNet’s convolution base with ImageNet weights instead of random values. The results of the proposed model under intra-sensor, same-material, and cross-material protocols are discussed in Section 4.
@@ 4.1 Database
Experiments are conducted on the LivDet 2015, 2017, and 2019 databases to validate the performance of the proposed model. Each database includes multiple sensing devices, with fingerprint samples from these devices organized into separate datasets for training and testing. The details of utilized databases including the name of sensing devices, number of training and testing samples, and names of the fabrication materials are mentioned in Table I.
TABLE I: Details of the databases utilized for the validation of DyFFPAD
@@
LivDet 2015 [23] Crossmatch 1473/1448 1000/1500 Ecoflex, Body Double, Playdoh, OOMOO, Gelatine
@@
LivDet 2017 [38] Greenbit 1200/2040 1000/1700 Gelatine, Latex, Liquid Ecoflex, Body Double, Ecoflex, Woodglue
@@
LivDet 2019 [25] Greenbit 1200/1224 1000/1020 Woodglue, Mix1, Mix2, Liquid Ecoflex , Body Double, Ecoflex
@@ 4.2 Performance Metrics
The performance of the proposed model is assessed using the metrics specified by ISO/IEC IS 30107 [1]. The Attack Presentation Classification Error Rate (APCER) measures the proportion of spoof fingerprint samples that are misclassified. The proportion of real fingerprint samples that are misclassified as spoofs is measured by the Bonafide Presentation Classification Error Rate (BPCER). Eq. (4) and Eq. (5) denote APCER and BPCER, respectively.
$APCER=\frac{\textnormal{\# Mis-classified spoof samples}}{\textnormal{Total fake samples}}\times{100}$ (4)
$BPCER=\frac{\textnormal{\# Mis-classified live samples}}{\textnormal{Total live samples}}\times{100}$ (5)
The Average Classification Error (ACE) is determined by averaging APCER and BPCER and is used to assess the overall performance of the FPAD system. The formulation of ACE is represented by Eq. (6).
@@
The ACE is subsequently used to determine the proposed model’s accuracy which is formulated as Eq. (7).
@@ 4.3 Implementation Details
The proposed model is implemented in Python and utilizes the TensorFlow-Keras library. The LPQ and handcrafted features are extracted using Matlab. The NVIDIA TESLA P100 GPU was used for all training and testing. Each model is trained for 300 epochs which requires 4 to 5 hours for its training.
@@ 4.4 Ablation Study
Figure 5: Receiver Operating Characteristic (ROC) curves for LivDet 2017 Orcanthus, Digital Persona, and Greenbit sensors
In order to compare the performance of the proposed model with handcrafted features, LPQ image descriptor, and DenseNet classifier, we have trained the corresponding DNN and CNN models. The handcrafted features and LPQ obtained from the fingerprint samples are used for the training of DNN while DenseNet-121 is trained on original fingerprint samples. The performance of the individual models and the proposed model is validated using LivDet 2017 which is a challenging database since it has training and testing sets containing spoof samples from different materials. The findings of this experiment are reported in Table II. Table II clearly indicates the supremacy of DyFFPAD over the DNN trained with LPQ, DNN trained with handcrafted features and LPQ, and the DenseNet classifier. The dynamic ensemble of handcrafted and deep features empowers the model to attain better classification accuracy irrespective of the materials used for the fabrication of spoofs. The proposed DyFFPAD models attain an overall classification accuracy of 96.55% as compared with 79.12% (LPQ+DNN), 75.93% (Handcrafted+DNN), 90.91% (LPQ+Handcrafted+DNN) and 93.64% (DenseNet-121 classifier). The performance comparison of the aforementioned models with DyFFPAD is also shown using the Receiver Operating Characteristics (ROC) curve which is depicted in Fig. 5. ROC curve is a representation of variation in True Positive Rate (TPR) and False Positive Rate (FPR) and is essential in measuring the performance of a machine/deep learning model. The Area Under the Curve (AUC) represents the region enclosed by the plot of the FPR versus TPR. A larger AUC indicates the better proficiency of the model while classifying the input samples.
5 Experimental Results and Comparative Analysis
5.1 Experimental Results
The performance of the proposed method has been validated in two benchmark scenarios: intra-sensor and known spoofing material, and intra-sensor and unknown spoofing material. The following subsections provide detailed descriptions of these scenarios and the corresponding results of the proposed method.
TABLE II: Findings of the ablation study performed on LivDet 2017 database
@@ 5.1 Experimental Results
Accuracy
Accuracy
(Digital Persona)
Accuracy
@@
TABLE III: Performance on LivDet 2015 database
Database Sensor BPCER
@@
5.1.1 Intra-Sensor and Known Spoofing Material
The training and testing fingerprint samples are acquired using the same sensing device in this experimental setup. Moreover, the spoof samples in training and testing datasets are fabricated using the same materials. LivDet 2015 partially follows this protocol as it consists of two-thirds of the testing spoofs fabricated with known materials and one-third with unknown materials. The results of the proposed model on the LivDet 2015 database are presented in Table III. The table indicates that the proposed model achieves an average BPCER of 5.79% and an APCER of 2.28%, as shown in the column labeled ”APCER (Known).”
TABLE IV: Performance on LivDet 2017 and 2019 databases
Database Sensor BPCER APCER ACE (%) EER (%)
@@ 5.1.1 Intra-Sensor and Known Spoofing Material
5.1.2 Intra-Sensor and Unknown Spoofing Material
In this experimental protocol, the training and testing fingerprint samples are acquired using the same sensing device. However, the materials used for the fabrication of training and testing spoof samples are different. This validation protocol is designed to evaluate the FPAD system’s efficacy in a real-attack scenario, wherein an adversary may attempt to present a forged fingerprint artifact created with novel fabrication materials that were not part of the FPAD model’s training data. LivDet 2017 and 2019 are prepared as per this protocol of FPAD. The results of our proposed method, as evaluated on the databases mentioned, are summarized in Table IV. As depicted in Table IV, our proposed model achieves a BPCER of 4.53%, APCER of 2.64%, and an ACE of 3.5%, and EER of 3.6% on the LivDet 2017 database. Similarly, on the LivDet 2019 database, the proposed model showcases the ability to discriminate between live and spoof samples, with a BPCER of 5.91%, APCER of 2.39%, and EER of 4.23%. Notably, when confronted with the spoof samples contained in the LivDet 2015 database, the proposed method exhibits an average APCER of 3.28%, as indicated in the column labeled ”APCER (unknown)” in Table III. It also shows that the proposed methods achieve an ACE of 3.90% and EER of 4.57% on LivDet 2015 database. The live and spoof fingerprint samples exhibit different textures and ridge valley widths due to significant differences between the properties of finger skin and spoofing materials. The convolutional layers of CNNs allow for effective classification by extracting distinguishing features from spoof samples, regardless of the materials used for their fabrication. Moreover, the handcrafted features and image descriptor extract discriminating features for the training of DNN. The dynamic ensemble of DNN and CNN facilitates improved parameter learning for both networks. The proposed architecture reduces the need for training two different models. The findings of the proposed method are compared to existing methods validated on benchmark databases, as discussed in section 5.2.

### v5

@@ 3 Proposed Work
In this paper, an ensemble of deep and handcrafted features is proposed for the detection of spoofs in different evaluation protocols of FPAD. It consists of two modules that collaborate to detect the PAs. The first module is a DNN that is fed with the Local Phase Quantization (LPQ) features and existing handcrafted features. The second module is the DenseNet CNN classifier which has shown remarkable performance while being validated on various image databases including imagenet Deng et al. (2009). Both of the modules are fused dynamically which results in better learning of them as compared with their individual performances. The details of LPQ, handcrafted features, DenseNet, and DyFFPAD are provided in the following sub-sections.
@@ 3.1 Local Phase Quantization (LPQ)
LPQ Ghiani et al. (2012) works on the Fourier transform’s blur-insensitive property which makes it robust against the redundant information and blur present in the input sample. We have utilized it as a prominent feature due to its capability of exploiting the minute information that gets missing in the fabrication process of a spoof. The formulation of the LPQ descriptor is denoted as Eq. 1.
@@ 3.2 Handcrafted Features
The fabrication process impacts the quality of the spoofs fabricated with various materials. According to recent research Sharma and Dey (2019) Sharma and Selwal (2021), handcrafted features are proven to be useful while detecting PAs. In this work, we have utilized some existing handcrafted features Sharma and Dey (2019), Šmida et al. (2015) including ridge-valley clarity, gabor quality, ridge-valley smoothness, number of abnormal ridges and valleys, orientation certainty level, and frequency domain analysis to estimate the quality of the input fingerprint sample. The details of these features is provided in the following subsections.
@@ 3.2.1 Ridge and Valley Smoothness (RVS and VWS)
It denotes the smoothness of ridge width and valley width which is exhibited by the live sample but not by the spoof. This irregularity in the spoofs is caused because of the varying elasticity levels of spoof fabrication materials. These materials do not possess the elasticity same as human skin. In addition to this, the pressure applied on the fingertip at the time of sample collection is also a potential reason behind this. RWS and VWS are computed block-wise by cropping the vertical ridge-valley structure after the rotation and binarization. Thereafter, pixels of the block are labeled as ridge or valley with the help of a linear regression algorithm. For each horizontal line of the block, ridge and valley width is computed. RWS and VWS features are obtained by averaging the standard deviation of ridge and valley widths.
@@ 3.2.2 Frequency Domain Analysis (FDA)
FDA of a local block is extracted by computing the ridge-valley structure’s 1D signature. The frequency of the sinusoidal ridge-valley structure is calculated using the discrete Fourier transform of this 1D signature. FDA is consistent in live fingerprints while it is having a varying frequency of sinusoidal ridge-valley patterns in spoof fingerprints.
@@ 3.2.4 Gabor Quality (G)
To calculate gabor quality at the local level, filter bank with many orientation on each pixel is applied. For a small number of filters with similar orientations as of block, the Gabor response with a normal ridge-valley pattern remains high, but it is low and steady for the same block with an improper ridge-valley structure. Lastly, the standard deviation of the gabor filter bank’s output is used to determine the Gabor quality (G) of a block. Gabor filter is well-suited for texture analysis, feature extraction, edge detection, etc.
@@ 3.2.5 Ridge-Valley Clarity (RVC)
The ridge and valleys present in live fingerprints block are found to be separated consistently. This happens due to the uniform elasticity property of human finger skin.This separation, on the other hand, differs in spoof samples due to inappropriate widths of ridges and valleys induced by the different elasticity levels of spoofing materials employed in spoof fabrication. The RVC feature is computed by dividing the original sample into fixed-sized blocks and then computing the average widths of ridge and valley in the block.
3.2.6 Abnormal ridges and valleys ($R_{ab}$ and $V_{ab}$)
The width of ridges and valleys is found to be in a range of 5 to 10 pixels in samples of human fingers captured from a sensing device with a resolution of 500 dpi. Since the human skin and spoofing materials have different levels of elasticity, the width of ridges and valleys differs in corresponding samples. An abnormality of ridge or valley is identified if its width divergence exceeds a particular threshold in various rows of the block.
@@ 3.2.7 Feature vector from quality features
The final feature vector representing the collection of all above discussed features is represented as Eq. (2) is generated using the mean ($\mu$) and standard deviation ($\sigma$) of the features discussed above.
@@ 3.3 DenseNet-121
DenseNetHuang et al. (2017) consists of convolutional layers connected to their previous layers in a feed-forward manner. The majority of CNN designs suffer from the vanishing gradient problem, which occurs when the Rectified Linear Unit (ReLU) is used as an activation function. In deeper CNNs, the gradient of the loss function approaches zero due to the use of ReLU activation functions. This phenomenon further impacts the learning process which results in overfitting of the model. As fingerprint images have limited color and texture information, typical CNNs encounter this issue when they are utilized in this domain. This problem of vanishing gradient is solved by DenseNet which has skip connections among the layers inside convolutional blocks as shown in Fig. 2.
@@
DenseNet-121 comprises four dense blocks containing six, twelve, twenty-four, and sixteen convolutional layers, respectively. The architecture of DenseNet is illustrated in Figure 2. As per this architecture, each dense block performs three operations: batch normalization, ReLU activation and convolution, and aggregates feature maps from all its preceding layers. The output of a dense block is formulated as shown in Eq. (3).
@@
where $F_{n}$ denotes a function that performs batch normalization followed by a convolution operation and $A_{0},A_{1},A_{n-1}$ are representing the concatenation of all feature maps from the previous layers. In addition, after each dense block, a transition block is utilized to execute a convolution operation with a kernel size of $1\times 1$. Following this convolution procedure, a pooling operator minimizes the size of feature maps after each dense block. Figure 3 depicts the internal architecture of the dense layer and transition layer.
@@ 3.4 Deep Neural Network (DNN)
DNN is an artificial neural network that has several fully connected layers. The working of DNNs is inspired by the human brain as its neurons resemble similar characteristics as neurons of the brain. We have used DNN to test the individual performance of the handcrafted features and LPQ image descriptor and to design the proposed architecture.
@@ 3.5 Proposed DyFFPAD Model
The architecture of the proposed DyFFPAD is depicted in Fig. 4. It is a dynamic ensemble of handcrafted and convolutional features. The DNN works on the feature vectors generated from handcrafted features and LPQ image descriptors while the convolution base of DenseNet extracts convolutional features from the input samples. The DNN 1 in Fig. 4 which consists of three dense layers having 256, 128, and 32, neurons, respectively is fed with the features extracted from LPQ image descriptor and handcrafted features. Similarly, DNN 2 is composed of three dense layers having 1024, 256, and 32, neurons, respectively. Finally, the outcome of both DNNs is combined to form a feature vector which is further fed as an input to another DNN 3 that consists of fully connected layers with 48, 16, and 1 neurons, respectively. In the forward pass of the training process, a confidence score is produced by the model for an input sample. The loss of the output score with respect to the expected output is then calculated which is back-propagated to the entire model including DenseNet as well as DNNs for the learning of their parameters. The proposed model works in an end-to-end manner for the forward and backward pass of the training process and reduces the limitations of the static ensemble of handcrafted features and convolutional features for the accomplishment of a classification task.
@@ 3.5.1 Pre-processing of input samples
The fingerprint samples captured with the sensing devices have white space around the fingertip impression. This white space is meaningless and is required to be removed for the extraction of the desired features from fingerprint samples. Therefore, a region of interest from the input fingerprint sample is identified as in pre-processing.
@@ 3.5.3 Training of DyFFPAD
The proposed DyFFPAD model is trained on benchmark LivDet databases from scratch. To expedite the training process, we initialized the parameters of DenseNet’s convolution base with ImageNet weights instead of random values. The results of the proposed model under intra-sensor, same-material, and cross-material protocols are discussed in Section 4.
@@ 4.1 Database
The proposed work is evaluated on the LivDet 2015, 2017, and 2019 databases to assess the performance widely used publicly available database. Each database includes multiple sensing devices, with fingerprint samples from these devices organized into separate datasets for training and testing. The details of utilized databases including the name of sensing devices, number of samples in training and testing, and names of the fabrication materials are mentioned in Table 1.
Table 1: Details of the databases utilized for the validation of DyFFPAD
@@
LivDet 2015 Mura et al. (2015) Crossmatch 1473/1448 1000/1500 Ecoflex, Playdoh, Body Double, OOMOO, Gelatine
@@
LivDet 2017 Yambay et al. (2018) Greenbit 1200/2040 1000/1700 Gelatine, Latex, Liquid Ecoflex, Body Double, Ecoflex, Woodglue
@@
LivDet 2019 Orrù et al. (2019) Greenbit 1200/1224 1000/1020 Woodglue, Mix1, Mix2, Liquid Ecoflex , Body Double, Ecoflex
@@ 4.2 Performance Metrics
The metrics specified by ISO/IEC IS 30107 30107-3:2017 (en) are used to assess the performance of the proposed model. The performance is measures using Attack Presentation Classification Error Rate (APCER) and Bonafide Presentation Classification Error Rate (BPCER) metrics. APCER measures the proportion of spoof fingerprint samples that are misclassified. The proportion of real fingerprint samples that are misclassified as spoofs is measured by BPCER. Eq. (4) and Eq. (5) denote APCER and BPCER, respectively.
$APCER=\frac{\textnormal{\# Mis-classified spoof samples}}{\textnormal{\# fake samples}}\times{100}$ (4)
$BPCER=\frac{\textnormal{\# Mis-classified live samples}}{\textnormal{\# live samples}}\times{100}$ (5)
The overall performance is evaluated using Average Classification Error (ACE) which is determined by averaging APCER and BPCER. The formulation of ACE is represented by Eq. (6).
@@
The ACE is also used to determine the proposed model’s accuracy which is formulated as Eq. (7).
@@ 4.3 Implementation Details
The proposed model is implemented using the TensorFlow-Keras library in Python. The NVIDIA TESLA P100 GPU was used for all training and testing. Each model is trained for 300 epochs which requires 4 to 5 hours for its training. The learning rate and batch size are kept as 0.0001 and 8, respectively.
@@ 4.4 Ablation Study
Figure 5: Receiver Operating Characteristic (ROC) curves for LivDet 2017 datasets (Orcanthus, Digital Persona, and Greenbit)
To assess, the performance of the proposed model with handcrafted features, LPQ image descriptor, and DenseNet classifier, we have trained the corresponding DNN and CNN models. The handcrafted features and LPQ obtained from the fingerprint samples are used for the training of DNN while DenseNet-121 is trained on original fingerprint samples. The performance of the individual models and the proposed model is validated using LivDet 2017 which is a challenging database since it contains spoof samples of different materials in training and testing sets. The results of this experiment are presented in Table 2 which clearly indicates the supremacy of DyFFPAD over the DNN trained with LPQ, DNN trained with handcrafted features and LPQ, and the DenseNet classifier. The dynamic ensemble of handcrafted and deep features empowers the model to attain better classification accuracy irrespective of the materials used for the fabrication of spoofs. The proposed DyFFPAD models attain an overall classification accuracy of 96.55% as compared with 79.12% (LPQ+DNN), 75.93% (Handcrafted+DNN), 90.91% (LPQ+Handcrafted+DNN) and 93.64% (DenseNet-121 classifier). The performance comparison of the aforementioned models with DyFFPAD is also shown using the Receiver Operating Characteristics (ROC) curve which is depicted in Fig. 5. ROC curve represents relationship between True Positive Rate (TPR) and False Positive Rate (FPR) and is essential in measuring the performance of a machine/deep learning model. The Area Under the Curve (AUC) represents the region enclosed by the plot of the FPR versus TPR. A larger AUC indicates the better proficiency of the model while classifying the input samples.
5 Experimental Results
The performance has been validated in two benchmark scenarios: intra-sensor (known spoofing material), and intra-sensor (unknown spoofing material). The following subsections provide detailed descriptions and results of these evaluation protocols.
Table 2: Findings of the ablation study performed on LivDet 2017 database in terms of accuracy
@@ 5 Experimental Results
Orcanthus
Digital Persona
Greenbit
@@
Table 3: Experimental results of our proposed method on LivDet 2015 database
Database Sensor
BPCER
@@
5.0.1 Intra-Sensor (Known Spoofing Material)
The samples of the training and testing sets are acquired using the same sensing device and they are fabricated using the same materials in both the sets. LivDet 2015 partially follows this protocol as it consists of two-thirds of the testing spoofs fabricated with known materials and one-third with unknown materials. The results on the LivDet 2015 database are presented in Table 3. The results presented in table shows that the proposed model achieves an average APCER (known) of 2.28% and BPCER of 5.79%.
Table 4: Experimental results of our proposed method on LivDet 2017 and 2019 databases
Database Sensor
BPCER
APCER
ACE
EER
@@ 5.0.1 Intra-Sensor (Known Spoofing Material)
5.0.2 Intra-Sensor (Unknown Spoofing Material)
In this experimental protocol, the samples of the training and testing fingerprint samples are acquired using the same sensing device but the fabrication material used is different in both sets. This validation protocol is designed to evaluate the FPAD system’s efficacy in a real-attack scenario, wherein an adversary may attempt to present a forged fingerprint artifact created with novel fabrication materials that were not part of the FPAD model’s training data. LivDet 2017 and 2019 are prepared as per this protocol of FPAD. The results of this evaluation protocol are presented in Table 4. As given in Table 4, our model achieves a BPCER of 4.53%, APCER of 2.64%, and an ACE of 3.5%, and EER of 3.6% on the LivDet 2017 database. Similarly, on the LivDet 2019 database, the proposed model showcases the ability to differentiate between live and spoof samples, with a BPCER of 5.91%, APCER of 2.39%, and EER of 4.23%. Notably, when confronted with the spoof samples contained in the LivDet 2015 database, the proposed method exhibits an average APCER (unknown) of 3.28%, as indicated in Table 3. It also shows that the proposed methods achieve an ACE of 3.90% and EER of 4.57% on LivDet 2015 database. The live and spoof fingerprint samples exhibit different textures and ridge valley widths due to significant differences between the properties of finger skin and spoofing materials. The convolutional layers of CNNs allow for effective classification by extracting distinguishing features from spoof samples, regardless of the materials used for their fabrication. Moreover, the handcrafted features and image descriptor extract discriminating features for the training of DNN. The dynamic ensemble of DNN and CNN facilitates improved parameter learning for both networks. The proposed architecture reduces the need for training two different models. A comparative study of the obtained results of our proposed method on benchmark databases is provided in section 5.1.